# Jose Tupayachi

### v1

@@ Abstract
Transportation remains a major contributor to greenhouse gas emissions, highlighting the urgency of transitioning toward sustainable alternatives such as electric vehicles (EVs). Yet, uneven spatial distribution and irregular utilization of charging infrastructure create challenges for both power grid stability and investment planning. This study introduces TW-GCN, a spatio-temporal forecasting framework that combines Graph Convolutional Networks with temporal architectures to predict EV charging demand in Tennessee, United States (U.S.). We utilize real-world traffic flows, weather conditions, and proprietary data provided by one of the largest EV infrastructure company in the U.S. to capture both spatial dependencies and temporal dynamics. Extensive experiments across varying lag horizons, clustering strategies, and sequence lengths reveal that mid-horizon (3-hour) forecasts achieve the best balance between responsiveness and stability, with 1DCNN consistently outperforming other temporal models. Regional analysis shows disparities in predictive accuracy across East, Middle, and West Tennessee, reflecting how station density, population, and local demand variability shape model performance. The proposed TW-GCN framework advances the integration of data-driven intelligence into EV infrastructure planning, supporting both sustainable mobility transitions and resilient grid management.
@@ 1 Introduction
Transportation is an indispensable part of human life, and with the exponential growth of the urban population, its importance continues to increase. In 2023 the United States (U.S.), transportation services contributed 8.9% to the gross domestic product (GDP). Adjusted for inflation, the demand for transportation rose by 5.5% in 2023, Despite its economic benefits, transportation also imposes several negative externalities, including traffic congestion and environmental pollution.
@@
Greenhouse gas emissions (CHG), a key driver of climate change and global warming, are primarily caused by the consumption of carbon-based fuels (e.g., gasoline) during energy use. In 2022, the transportation sector was reported as one of the two driving contributors in total CO2 emissions (Raimi et al., 2022). The transportation sector accounted for approximately 28% of total U.S. GHG emissions in 2022 (U.S. Environmental Protection Agency, 2025). Followed, in 2023, about 37% of total U.S. energy consumption was attributed to the transportation sector (U.S. Energy Information Administration, 2024a). Nevertheless, electricity accounted for roughly 37% of the energy used in transportation (U.S. Energy Information Administration, 2024b).
@@
EVs have emerged as a widely adopted alternative transportation method in addition to shared mobility, playing an important role in the decarbonization and electrification of the transportation network (Cui et al., 2022). In the U.S. there were 196,000 publicly available EV stations reported in January 2025 (Climate Central, 2025). Further, the Bipartisan Infrastructure Law led by the U.S. administration investedFederal Highway Administration (2025) \$2.5 billion through the Charging and Fueling Infrastructure Discretionary Grant Program and the National Electric Vehicle Infrastructure Formula Program, including the most recent \$635 million in grants to expand publicly EV charging across the country.
@@
Nevertheless, the growing adoption of EVs results in an imbalance between supply and demand, especially at commercial charging stations. In certain areas, EV stations are heavily used, resulting in excessive pressure on power grids (Roy et al., 2023). In contrast, under-utilization in other regions can cause inefficient resource allocation and financial losses for investors and operators (Etxandi-Santolaya et al., 2023). Therefore, understanding the behavior of EV charging demand is crucial. This involves not only analyzing the spatial distribution of EV stations but also considering factors such as temporal features, weather-related features, urban geographical features, and traffic-based data.
@@
Since 2010, research on EV demand prediction has gained increasing attention, addressing both load and usage perspectives. To tackle this problem, scholars have explored a range of approaches, including probabilistic models (Tang and Wang, 2015), time series analysis (Amini et al., 2016), machine learning models (Liu et al., 2018), and deep learning techniques (Wang et al., 2023b). These demand predictions can be categorized into two groups: (a) long-term demand prediction (e.g., yearly) and (b) short-term demand prediction (e.g., hourly). Short-term predictions are particularly beneficial for dynamic price adjustments and managing load demands, while long-term predictions aid in infrastructure development and urban planning.
@@
To the best of our knowledge, no prior study has analyzed real-world U.S.-based (i.e., Tennessee) EV charging demand data while simultaneously accounting for spatial, weather-related, and traffic-based factors. In this study, we focus on Tennessee, a key state in the Southeastern U.S., where detailed datasets of charging station activity, traffic flow, and weather conditions are available. Using these real-world operational datasets, we aim to forecast energy consumption (in kWh) at EV charging stations. To achieve this, we develop a Traffic-Weather Graph Convolutional Network (TW-GCN) that integrates historical traffic and weather time series into a stacked GCN architecture. The model captures both temporal and spatial dependencies in graph-structured data to predict EV charging demand. Additionally, our multi-dimensional approach identifies clusters of stations with similar usage patterns, deepening insights into the spatial and temporal dynamics shaping EV charging demand in the Southeastern U.S. In this regard, our TW-GCN stretches existing approaches by integrating multimodal real-world data with spatio-temporal learning to deliver region-specific, accurate EV charging demand forecasts that can inform both infrastructure planning and grid resilience.
@@
The paper is organized as follows: Section 2 reviews existing work on EV station demand prediction and deep learning methods. Section 3 describes the datasets, preprocessing, and quality checks. Section 4 outlines the model architecture, training, and evaluation metrics. Section 5 presents results, model comparisons, and statistical analyses. Section 6 focuses on a Tennessee case study across short-, mid-, and long-term horizons. Finally, Section 7 summarizes findings and future research directions.
@@ 2 Literature Review
In this section, we review the existing studies from two perspectives. First, we discuss the evolution of Spatial-Temporal Graph Neural Networks and our reasoning for choosing them as our solution technique (see Section 2.1). Next, we examine state-of-the-art methods for EV charging demand prediction (see Section 2.2), where the increasing availability of historical data has led many researchers to employ deep learning approaches. Finally, we outline our contributions in Section 2.3.
@@ 2.1 Emergence of Spatial-temporal Graph Neural Networks & Graph Convolutional Networks
In recent years, the increasing availability of spatial-temporal data has led to the development of advanced deep learning models capable of capturing both spatial dependencies and temporal dynamics. Convolutional Neural Networks (CNNs) have demonstrated considerable efficiency, particularly when handling grid-like data structures such as images and spreadsheets (Li et al., 2021). However, many real-world data structures are irregular, including but not limited to protein-protein interaction networks, transportation networks, and infrastructure networks. To address this issue, Recurrent Neural Networks (RNNs) have been employed by many researchers (Salehinejad et al., 2017). Yet, they face certain challenges, such as vanishing and exploding gradients, difficulties in learning long-term dependencies, and restrictions to sequential data processing, such as time series. To overcome these limitations, GNNs have been proposed to extend convolutions to graph data structures (see (Wu et al., 2020) for an extensive review). Initial studies in this field predominantly focused on spectral representations based on the eigenvalues and eigenvectors of the graph Laplacian matrix (i.e., L=ADL=A-D where $A$ and $D$ are adjacency and a diagonal matrix where each entry $D_{ii}$ is the degree of node $i$, respectively) to perform graph convolutions (Bruna et al., 2013). Due to the computational burden associated with these spectral methods, research shifted towards spatially localized filters (Berg et al., 2017).
@@
However, as stated by Velickovic et al. (2017), the learned filters from these methods do not generalize well to different graph structures in terms of node and edge connections. This limitation led to the development of non-spectral approaches such as GraphSAGE, which creates node embeddings through sampling and aggregating features from local neighborhoods (Hamilton et al., 2017). A significant advancement was made by the introduction of Graph Attention Networks (GAT) Velickovic et al. (2017), which combine self-attention mechanisms with convolutional methods to address the node classification problem. Incorporating spatial-temporal dimensions into GNNs has further enhanced their applicability, particularly in dynamic networks where nodes and edges change over time. Spatial-temporal GNNs integrate spatial dependencies between nodes with temporal evolution (Yu et al., 2017). ST-GNNs have been successfully employed in traffic flow forecasting (Ali et al., 2022), traffic accident forecasting (Yu et al., 2021), EV demand prediction (Wang et al., 2023a), and passenger demand prediction (Tang et al., 2021). These models are particularly well suited for EV charging demand forecasting because this problem is inherently spatial-temporal. EV charging demand is influenced not only by time-dependent usage cycles (e.g., commute times, weekdays versus weekends, seasonal variations) but also by spatial correlations between stations that arise from road connectivity, geographic clustering of demand, and heterogeneous adoption patterns. Conventional forecasting models, such as regression, time series analysis, or deep learning architectures, often treat space and time independently, which limits their ability to capture cross-location spillover effects and evolving network interactions. In contrast, ST-GNNs are explicitly designed to learn from both spatial topologies and temporal dynamics, making them more effective in modeling when and where demand will arise. Building on these strengths, we propose a novel approach to forecast the demand for electric vehicle charging in the Southeastern U.S. (e.g., Tennessee region).
@@ 2.2 EV Charging Demand Forecasting
The interest in EV charging station demand forecasting has been growing exponentially. Early studies primarily focused on EV charging station load, which involves forecasting the electrical load measured in kilowatts or megawatts (Qian et al., 2010). One of the early studies was conducted by (Xie et al., 2011), where the authors utilized neural networks (NNs), radial basis function (RBF) NNs, and a prediction model to estimate the daily load of an EV charging station using historical data obtained from the Beijing Olympic Games EV Charging Station in 2010. Following this, (Xydas et al., 2013) analyzed a fleet of EVs to estimate the weekly charging demand in kW using decision tables, representative trees, NNs, and support vector machines (SVM). While these methods were effective in capturing temporal patterns of charging load, they often do not explicitly model spatial dependencies or interactions between different stations and external factors such as traffic or weather. To address this limitation, we adopt graph-based approaches that naturally incorporate both temporal and spatial relationships, motivating our choice of GNNs for the forecasting task.
@@
In contrast, Majidpour et al. Majidpour et al. (2014) utilized a time series method along with Random Forest (RF), SVM, and k-Nearest Neighbor (kNN) to estimate energy consumption at the charging outlet level. However, the authors did not use any spatial features or driving habit-based data. The following year, Olivella-Rosell et al. Olivella-Rosell et al. (2015) proposed an agent-based simulation model to understand the cumulative EV charging demand and its grid impact. One of the very first studies analyzed social (e.g., age) and economic (e.g., income) variables, which was further extended (Arias and Bae, 2016) with where real-world traffic and weather data were used to forecast EV charging demand; nevertheless, clustering and decision tree approach does not capture complex interactions between traffic, weather, and charging demand, particularly in dynamic, or under high EV penetration scenarios.
@@
Spatial-temporal charging demand prediction was conventionally tackled by simulation models and probabilistic models (Zhou and Lin, 2012; Zhang et al., 2015). Then extended by using the dynamic probabilistic models (Xia et al., 2019) where they considered driving behavior, battery usage, and traffic information. Moving forward, a short-term load forecast model (McBee et al., 2020) using SVM on European Commissions’ off Transport dataset where they considered the time of arrival, the departure time, the daily travel distance, and battery state of charge to forecast charging load. They also compared the proposed method with the Monte Carlo forecasting technique but did not account for critical factors like geographical location, weather, or traffic data. To address these limitations, Kim and Kim Kim and Kim (2021) introduced an ensemble machine learning model including ARIMA, NNs, and Long Short-Term Memory (LSTM) for forecasting EV charging demand. They aimed to enhance accuracy by considering factors such as charging patterns, weather conditions, and the day of the week. Further, the authors analyze three geographical scales station-, city-, and country-based.
@@
Other approaches have contemplated fuzzy sets, which are based on fixed-value rules, cannot be optimized through learning from big data.Among these, Zamee et al. Zamee et al. (2023) performed a MIC correlation to identify dominant features in asynchronously sampled datasets and applied a General Regression NN-based forecasting model. They selected dominant lagged exogenous variables through correlation multicollinearity analysis. Although they claim to handle big data and feature selection by correlation, the sources of the big data, the pipeline, and the reasons why certain features were considered significant have not been discussed. To develop a predictable and economically interpretable model (Kuang et al., 2024) a learning approach for accurate EV charging demand prediction and reasonable pricing, called PIAST integrating convolutional feature engineering, spatio-temporal dual attention mechanism and physics-informed neural network training (Raissi et al., 2019). They used the concept of price elasticity of demand (PED) which is a measure of how responsive the demand is to changes in its price. However, they ignored the impact of weather, traffic density, etc.
@@
Sun et al. Sun et al. (2021) investigated the role of EVs in supporting the smart grid through vehicle-to-grid (V2G) interactions. Using a one-month GPS trajectory dataset of 967 rental EVs in Beijing (January 2019), they applied fixed cutoff values on adjacency matrices to analyze charging and discharging patterns. However, this static analytical approach struggled to capture evolving dynamic demand patterns. Although the study highlights the value of examining demand from rental EVs, since they contribute substantially to the grid it overlooks how private and rental EVs together shape the overall load. Furthermore, relying solely on rental GPS trajectories limited the analysis: station-level demand variation due to geographical location, weather, or nearby amenities was not captured. Studying station-wise demand evolution could reveal how external factors drive charging behavior across regions. Building on this limitation, Yi et al. Yi et al. (2022) proposed a Seq2Seq model for forecasting monthly commercial EV charging demand. They complemented this with clustering techniques to capture regional spatiotemporal demand patterns. Yet clustering methods often treat data points in isolation, relying mainly on feature similarity. This restricts their ability to model richer contextual dependencies between locations. In contrast, GNNs can integrate multiple layers of context economic indicators, traffic flows, or environmental conditions, while explicitly modeling interlocation relationships.
@@
Recent studies have applied deep learning to EV charging demand forecasting, using approaches such as LSTMs for short-term station-level prediction (Wang et al., 2023b) and graph convolutional networks (GCNs) combined with gated recurrent units to capture spatio-temporal features (Wang et al., 2023a). GCNs, in particular, have demonstrated strong performance in modeling spatial relationships between charging stations and users (Hüttel et al., 2021; Wang et al., 2023c), making them well suited for capturing the dynamic and interconnected nature of EV demand. To further improve forecasting accuracy, it is important to incorporate external factors such as POI data, weather conditions, and traffic density, which can reveal evolving non-linear patterns in station-level demand. Table 1 summarizes these studies, highlighting their methodological innovations and key limitations that motivate our proposed approach.
Table 1: Summary of Selected EV Charging Demand Forecasting Studies
Authors Approach Contribution Limitations
Wang et al. Wang et al. (2023d) GCN (Spatio-Temporal) Models spatial-temporal EV demand in Beijing using learned dependencies Does not include traffic or weather data
Gunasekaran et al. Gunasekaran and Smith (2024) GCN–GRU Predicts charging using spatio-temporal GNNs Small U.S. dataset; no external features
Song et al. Song et al. (2023) AST-GIN Combines GCN + Informer using POIs and static weather features No dynamic external inputs; non-U.S. focus
Yan et al. Yan et al. (2020) Monte Carlo Simulation Simulates traffic and weather effects on EV load patterns Not predictive; only used for planning/simulation
Feng et al. Feng et al. (2023) Load Simulation Captures demand shifts due to traffic and temperature across zones Not used for real-time forecasting
Su et al. Su and Chow (2017) Traffic-Based Simulation Simulates charging demand based on traffic and routing data No machine learning; planning only

### v3

# Supervision, Writing – review & editing, Project Administration

@@ Abstract
Transportation remains a major contributor to greenhouse gas emissions, highlighting the urgency of transitioning toward sustainable alternatives such as Electric Vehicles (EVs). Yet, uneven spatial distribution and irregular utilization of charging infrastructure create challenges for both power grid stability and investment planning. This study introduces Traffic-Weather Graph Convolutional Network (TW-GCN), a spatio-temporal forecasting framework that combines Graph Convolutional Networks with temporal architectures to predict EV charging demand in Tennessee, United States. We utilize real-world traffic flows, weather conditions, and proprietary data provided by one of the largest U.S.-based EV infrastructure companies to capture both spatial dependencies and temporal dynamics. Extensive experiments across varying forecasting horizons, clustering strategies, and sequence lengths reveal that mid-horizon (3-hour) forecasts achieve the best balance between responsiveness and stability, with One-dimensional convolutional neural networks consistently outperforming other temporal models. Regional analysis shows disparities in predictive accuracy across East, Middle, and West Tennessee, reflecting how station density, Points of Interest and local demand variability shape model capabilities. The proposed TW-GCN framework advances the integration of data-driven intelligence into EV infrastructure planning while supporting sustainable mobility transitions.
@@ 1 Introduction
Transportation is an indispensable part of human life, and with the exponential growth of the urban population, its importance continues to increase. In 2023, transportation services in the United States (U.S.) contributed 8.9% to the gross domestic product. Adjusted for inflation, the demand for transportation increased by 5.5% in 2023. Despite its economic benefits, transportation also imposes several negative externalities, including traffic congestion and environmental pollution.
@@
Greenhouse gas (GHG) emissions, a key driver of climate change and global warming, are primarily caused by the consumption of carbon fuels (e.g., gasoline) during energy use. In 2022, the transportation sector was reported as one of the two main contributors to total CO2 emissions (raimi2022global). The transportation sector accounted for approximately 28% of total U.S. GHG emissions in 2022 (EPA2025). Subsequently, in 2023, the transportation sector accounted for about 37% of total U.S. energy consumption (EIA_US_Energy_Facts_2024).
@@
Electric vehicles (EVs), together with shared mobility, play a vital role in the decarbonization and electrification of transportation networks (cui2022battery). As of January 2025, the U.S. had approximately 196,000 publicly available EV charging stations (ClimateCentral2025_EVChargeUp). To support this transition, the Bipartisan Infrastructure Law allocated (FHWA2025evcharging) \$2.5 billion through the Discretionary Grant Program for Charging and Fueling Infrastructure and the National EV Infrastructure Formula Program, including a recent \$635 million in grants to further expand public EV charging nationwide.
@@
Nevertheless, the growing adoption of EVs results in an imbalance between supply and demand, especially at commercial charging stations. In certain regions, high utilization of EV charging stations exerts substantial stress on power grids (roy2023impact). In contrast, underutilization in other areas can cause inefficient resource allocation and financial losses for investors and operators (etxandi2023estimation). Therefore, understanding the behavior of EV charging demand is crucial. This involves not only analyzing the spatial distribution of EV stations but also considering factors such as temporal features, weather features, urban geographical features, and traffic-based data.
@@
To address this problem, scholars have explored a variety of approaches, such as probabilistic models (tang2015probabilistic), time series analyses (amini2016arima), machine learning algorithms (liu2018hybrid), and deep learning techniques (wang2023short). These demand prediction methods can be categorized into two groups: (a) long-term demand prediction (e.g., yearly) and (b) short-term demand prediction (e.g., hourly). Short-term predictions are particularly beneficial for dynamic price adjustments and managing load demands, while long-term predictions aid in infrastructure development and urban planning.
@@
To the best of our knowledge, no prior study has analyzed real-world U.S.-based (i.e., Tennessee) EV charging demand data while simultaneously accounting for spatial, weather, and traffic flows derived from TMCs metrics. In this study, we focus on Tennessee, a key state in the Southeastern U.S., where detailed datasets of charging station activity, traffic flow, and weather conditions are available. Using these real-world operational datasets, we aim to forecast energy consumption (in kWh) at EV charging stations. To achieve this, we propose a novel Traffic-Weather Graph Convolutional Network (TW-GCN) model that integrates historical traffic and weather time series within a stacked graph convolutional network (GCN) architecture. This model captures both temporal and spatial dependencies to predict the charging demand. Our multidimensional approach identifies clusters of stations with similar usage patterns to deliver region-specific, accurate EV charging demand forecasts that can inform both infrastructure planning and mobility transitions.
@@
The paper is organized as follows. Section 2 reviews existing work on EV station demand prediction and model architectures. Section 3 describes the datasets, preprocessing, and quality checks. Section 4 outlines the model architecture, training, and evaluation metrics. Section 5 presents results, model comparisons, and statistical analyses. Section 6 focuses on a Tennessee case study across short, mid, and long-term horizons. Finally, Section 7 presents the conclusions and outlines directions for future research.
@@ 2 Literature Review
In this section, we review the existing studies from two perspectives. First, we discuss the evolution of Spatial-temporal Graph Neural Networks and our reasoning for choosing them as our solution technique (Section 2.1). Next, we examine state-of-the-art methods for EV charging demand prediction (Section 2.2), where the increasing availability of historical data has led many researchers to employ deep learning approaches. Finally, we outline our contributions in Section 2.3.
@@ 2.1 Emergence of Spatial-temporal Graph Neural Networks & Graph Convolutional Networks
In recent years, the increasing availability of spatial-temporal data has led to the development of advanced deep learning models capable of capturing both spatial dependencies and temporal dynamics. Convolutional Neural Networks (CNNs) have demonstrated considerable efficiency, particularly when handling grid-like data structures such as images and spreadsheets (li2021survey). However, many real-world data structures are irregular, including but not limited to protein–protein interaction networks, transportation networks, and infrastructure networks. To address this issue, Recurrent Neural Networks (RNNs) have been employed by many researchers (salehinejad2017recent). Yet, they face certain challenges, such as vanishing and exploding gradients, difficulties in learning long-term dependencies, and restrictions related to sequential data processing, such as time series.
To overcome these limitations, Graph Neural Networks (GNNs) have been proposed to extend convolutions to graph data structures; see wu2020comprehensive for an extensive review. The term GNN serves as an umbrella concept encompassing a broad family of architectures designed to process graph-structured data. Within this family, several specialized models have been developed, such as GCNs, Graph Attention Networks (GATs), and Graph Recurrent Networks (GRNs), each introducing distinct mechanisms for information propagation and aggregation across nodes. Initial studies in this field predominantly focused on spectral representations based on the eigenvalues and eigenvectors of the graph Laplacian matrix (i.e., L=ADL=A-D, where $A$ is the adjacency matrix and $D$ is a diagonal matrix whose entries $D_{ii}$ represent the degree of node $i$) to perform graph convolutions (bruna2013spectral).
@@
Due to the computational burden associated with these spectral methods, research later shifted toward spatially localized filters (berg2017graph). However, as stated by velickovic2017graph, the learned filters from these methods do not generalize well to different graph structures in terms of node and edge connections. This limitation led to the development of non-spectral approaches such as GraphSAGE, which creates node embeddings through sampling and aggregating features from local neighborhoods (hamilton2017inductive). A significant advancement was made with the introduction of GATs by velickovic2017graph, which combine self-attention mechanisms with convolutional methods to address the node classification problem. Incorporating spatial-temporal dimensions into GNNs has further enhanced their applicability, particularly in dynamic networks where nodes and edges change over time. Spatial-temporal GNNs (ST-GNNs) integrate spatial dependencies between nodes with temporal evolution (yu2017spatio).
ST-GNNs have been successfully employed in traffic flow forecasting (ali2022exploiting), traffic accident forecasting (yu2021deep), demand prediction (wang2023predicting), and passenger demand prediction (tang2021multi). These models are particularly well-suited for charging demand forecasting since the problem is inherently spatial-temporal. EV charging demand is influenced not only by time-dependent usage cycles (e.g., commute times, weekdays versus weekends, seasonal variations), but also by spatial correlations between stations that arise from road connectivity, geographic clustering of demand, and heterogeneous adoption patterns. Conventional forecasting models such as regression, time series analysis, or deep learning architectures often treat space and time independently, which limits their ability to capture cross-location spillover effects and evolving network interactions. In contrast, ST-GNNs are explicitly designed to learn from both spatial topologies and temporal dynamics, making them more effective at modeling when and where demand will arise. Building on these strengths, we propose a novel approach to forecast the demand for EV charging in the Southeastern U.S. (i.e., the Tennessee region).
@@ 2.2 EV Charging Demand Forecasting
The interest in EV charging station demand forecasting has been growing exponentially. Early studies primarily focused on EV charging station load, which involves forecasting the electrical load measured in kilowatts or megawatts (qian2010load). In one of the earliest contributions, xie2011research applied neural networks (NNs), including radial basis function variants, alongside a predictive model to estimate daily EV charging station loads using data from the 2010 Beijing Olympic Games site. Following this, xydas2013electric analyzed a fleet of EVs to estimate the weekly charging demand in kW using decision tables, representative trees, NNs, and support vector machines (SVM). While these methods were effective in capturing temporal patterns of charging load, they often do not explicitly model spatial dependencies or interactions between different stations and external factors such as traffic or weather. To address this limitation, we adopt graph-based approaches that naturally incorporate both temporal and spatial relationships, motivating our choice of GNNs for the forecasting task.
@@
In contrast, majidpour2014novel utilized a time series method along with Random Forest (RF), SVM, and k-Nearest Neighbor (kNN) to estimate energy consumption at the charging outlet level. However, the authors did not use any spatial features or driving habit-based data. The following year, olivella2015probabilistic proposed an agent-based simulation model to understand the cumulative EV charging demand and its grid impact. One of the very first studies analyzed social (e.g., age) and economic (e.g., income) variables, which was further extended (arias2016electric) in which real-world traffic and weather data were used to forecast EV charging demand; nevertheless, clustering and decision tree approach does not capture complex interactions between traffic, weather, and charging demand, particularly in dynamic scenarios or under high EV penetration scenarios.
@@
Spatial-temporal charging demand prediction was conventionally tackled by simulation and probabilistic models (zhou2012spatial; zhang2015integrated), then extended by using dynamic probabilistic models (xia2019ev), where they considered driving behavior, battery usage, and traffic information. Moving forward, a short-term load forecasting model (mcbee2020forecasting) using SVM was developed based on the European Commission’s Off Transport dataset, where the time of arrival, departure time, daily travel distance, and battery state of charge were considered to forecast the charging load. The authors also compared the proposed method with the Monte Carlo forecasting technique but did not account for critical factors like geographical location, weather, or traffic data. To address these limitations, kim2021forecasting introduced an ensemble machine learning model including Autoregressive Integrated Moving Average (ARIMA), NNs, and Long Short-Term Memory (LSTM) for forecasting EV charging demand. They aimed to enhance accuracy by considering factors such as charging patterns, weather conditions, and the day of the week. Further, the authors analyze three geographical scales: station-, city-, and country-based.
@@
Other approaches have contemplated fuzzy sets, which are based on fixed-value rules, and cannot be optimized through learning from big data. Among these, zamee2023self performed a Maximal Information Coefficient correlation to identify dominant features in asynchronously sampled datasets and applied a General Regression NN-based forecasting model. They selected dominant lagged exogenous variables through correlation multi-collinearity analysis. Although they claim to handle big data and feature selection by correlation, the sources of the big data, the pipeline, and the reasons why certain features were considered significant have not been discussed. To develop a predictable and economically interpretable model, kuang2024physics proposed a learning approach for accurate EV charging demand prediction and rational pricing, called PIAST, which integrates convolutional feature engineering, a spatio-temporal dual attention mechanism, and physics-informed neural network training (raissi2019physics). They incorporated the concept of price elasticity of demand, which measures how responsive demand is to price changes. However, the model did not account for external factors such as weather and traffic density.
@@
sun2021exploring investigates the role of EVs in supporting the smart grid through vehicle-to-grid interactions. Using a one-month GPS trajectory dataset from 967 rental EVs in Beijing (January 2019), the authors applied fixed cutoff values to the adjacency matrices to analyze charging and discharging patterns. However, this static analytical approach failed to capture the evolving dynamics of demand patterns. Although the study highlights the importance of analyzing demand from rental EVs, which significantly contribute to the grid, it overlooks the combined impact of private and rental EVs on the overall load. Relying solely on rental GPS trajectories limited the analysis, as it failed to capture station-level demand variations influenced by geographic location, weather, or nearby amenities. Examining station demand evolution could reveal how such external factors shape charging behavior across regions.
Building on this limitation, yi2022electric proposed a Seq2Seq model for forecasting monthly commercial EV charging demand. The authors complemented this with clustering techniques to capture regional spatiotemporal demand patterns. Yet clustering methods often treat data points in isolation, relying mainly on feature similarity. This limitation hinders their ability to capture richer contextual dependencies between locations; in contrast, GNNs can incorporate multiple layers of context while explicitly modeling interlocation relationships.
@@
Recent studies have applied deep learning to EV charging demand forecasting, using approaches such as LSTMs for short-term station-level prediction (wang2023short) and GCNs combined with gated recurrent units to capture spatio-temporal features (wang2023predicting). GCNs, in particular, have demonstrated strong performance in modeling spatial relationships between charging stations and users (huttel2021forecasting; wang2023gcn) for capturing the dynamic and interconnected nature of EV demand. To further improve forecasting accuracy, it is important to incorporate external factors such as Points of Interest (POI) data, weather conditions, and traffic density, which can reveal evolving non-linear patterns in station-level demand. Table 1 summarizes these studies, highlighting their methodological innovations and key limitations that motivate our proposed approach.
Table 1: Recent Studies in Spatiotemporal Forecasting for EV Charging Demand
Authors Approach Contribution Limitations Coverage
wang2025adaptive Adaptive Spatial-Temporal Graph Recurrent Network (ASTGRN) Eliminates the need for predefined geographic distances or prior spatial knowledge Does not incorporate traffic inputs and evaluated only on GRU architecture Beijing, China
li2025spatiotemporal Non-negative Matrix Factorization (NMF) to extract latent spatiotemporal EV charging patterns from high-resolution driver data Large-scale driver data revealed three main charging patterns: late night, daytime, and early evening Excludes external factors such as weather, user behavior, and traffic Four major cities in China
cui2025weather Random Forest combined with SHAP values Identification of weather in origin–destination pairs to affect long-distance highway travel for EVs Model may not generalize to colder regions, high-traffic periods, or areas with different infrastructure Guangdong, China
huang2025energy Energy Consumption Transformer (ECT) Addresses EV data scarcity using a shot-based learning approach with transferable prior knowledge. Limited generalization to regions with distinct driving patterns or environmental conditions. Shanghai, China
chen2025disaggregation Bi-level Spatio-Temporal Graph Convolutional Network (ST-GCN) framework. Event and design dissagregation for non-intrusive community-level EV charging profile disaggregation Tested on limited GRU, Transformer, and LSTM architectures, with no further benchmark analysis Selected states in the United States
zheng2025spatio Spatio-Temporal Charging Load Prediction Method based on Cellular Traffic Simulation (CTM) Real-time traffic parameters alongside charging load predictions, allowing charging station planning and traffic-aware charging guidance Only analyses a single urban region Hunan, China
shahrokhi2025two Two-Stage Clustering and Classification Combines spatial (where) and temporal (when) charging behaviors to group EV users for accurate grid impact forecasting Primarily focuses on temporal features, with limited consideration of spatial mobility and accessibility factors Saskatchewan, Canada
wu2025spatio Spatio-temporal prediction framework for EV charging load based on real road networks Incorporates varied EV types and trip behavior to model energy consumption Limited to a single city and simulation-based validation Yunnan, China
tian2025multi Multi-Scale Spatial-Temporal Graph Attention Network (MSSTGAN) Uses graph attention-based gated fusion to dynamically integrate spatio-temporal embeddings, capturing the importance of different stations (spatial) and historical time points (temporal) Evaluated on geographically dispersed networks without traffic or weather data Palo Alto, Boulder (U.S.), Perth (U.K.)